OpenStudy (anonymous):

A rectangle has a width of 3x + 5 feet and a length of 20 feet. The perimeter is 67 feet. Find the constant that represents the width.

OpenStudy (anonymous):

2(3x+5) + 2(20) = 67 2(3x+5) = 27 6x = 17 x = 17/6

OpenStudy (selena2345):

hi tasnia it is me simran